OpenStudy (anonymous):

Which ordered pair is a solution of the equation y = x – 2? (1 point) (3, 1) (1, 3) (–1, 3) (3, –1)

OpenStudy (anonymous):

All you need to do is plug in the x values of each ordered pair and see which one yields a true statement. (3, -1) (option D) y = x - 2 -1 = 3 - 2 -1 = 1 Is this true? Of course not. So we can eliminate D. Do this for the others and you'll find your answer!

OpenStudy (anonymous):

Thank you very much c;